A rectangular loop of wire lays flat on a table in a uniform magnetic field of 4.0x10^-2 T. The length of the rectangle is 2.5x10^-2 m, and the width is stretched smoothly from 5.0x10^-2 m to 7.5x10^-2 m in 0.3 seconds. How much emf is induced in the loop?
Answer #1

Flux = magnetic field * area

emf = rate of change of flux

emf = B*l*Δw/t = 4*10-2* 2.5*10-2 * 2.5*10-2 / .3 = 8.33 * 10-5 V